OpenStudy (sunshinehaggard):

(1 + 2i)(2 + i) + = 5(2 + i)

OpenStudy (sunshinehaggard):

@kayders1997

OpenStudy (sunshinehaggard):

@Matt6288

OpenStudy (sunshinehaggard):

@Directrix

OpenStudy (sunshinehaggard):

@MathMusician

OpenStudy (kendricklamar2014):

Is this the equation:\[(1 + 2i)(2 + i) + \] @sunshinehaggard

OpenStudy (sunshinehaggard):

yes im trying to find the missing term to the equation

OpenStudy (sunshinehaggard):

@KendrickLamar2014

hartnn (hartnn):

let the missing term be x+iy (1 + 2i)(2 + i) + x+iy = 5(2 + i) now collect all the terms with and without 'i' on both sides compare real and imaginary parts :) try it , let me know if you get stuck.

OpenStudy (sunshinehaggard):

it was wrong

hartnn (hartnn):

can you show what have you done?

hartnn (hartnn):

(1 + 2i)(2 + i) + x+iy = 5(2 + i) 2 + 4i + i + 2i^2 + x + iy = 10 + 5i so, 2-2 +x = 10 x= 10 5i +iy = 5i so, y = 0 so, did you put 10 in the blank?

OpenStudy (sunshinehaggard):

i know where i messed up thank you

hartnn (hartnn):

welcome ^_^